# Decoding Futures Price Dynamics: A Regularized Sparse Autoencoder for Interpretable Multi-Horizon Forecasting and Factor Discovery

### v1

@@ Abstract
Commodity price volatility creates economic challenges, necessitating accurate multi-horizon forecasting. Predicting prices for commodities like copper and crude oil is complicated by diverse interacting factors (macroeconomic, supply/demand, geopolitical, etc.). Current models often lack transparency, limiting strategic use. This paper presents a Regularized Sparse Autoencoder (RSAE), a deep learning framework for simultaneous multi-horizon commodity price prediction and discovery of interpretable latent market drivers. The RSAE forecasts prices at multiple horizons (e.g., 1-day, 1-week, 1-month) using multivariate time series. Crucially, L1 regularization ($\|\mathbf{z}\|_{1}$) on its latent vector $\mathbf{z}$ enforces sparsity, promoting parsimonious explanations of market dynamics through learned factors representing underlying drivers (e.g., demand, supply shocks). Drawing from energy-based models and sparse coding, the RSAE optimizes predictive accuracy while learning sparse representations. Evaluated on historical Copper and Crude Oil futures data with numerous indicators, our findings indicate the RSAE offers competitive multi-horizon forecasting accuracy and data-driven insights into price dynamics via its interpretable latent space, a key advantage over traditional black-box approaches.
@@ 2.1 Introduction to Commodity Price Forecasting: Significance and Challenges
Accurate forecasting of commodity prices is crucial, given their extensive impact on economies and daily life [1]. Variations in prices of essential goods—from agricultural produce to energy resources and industrial metals—affect many aspects of economic activities and individual well-being. For instance, volatile prices can significantly influence agricultural productivity and farmer incomes [35], while also presenting unpredictable expenses for consumers. Countries heavily dependent on commodity trade find their macroeconomic stability closely linked to commodity price fluctuations [18]. Additionally, amid rising concerns about global food security, effective prediction of staple commodity prices becomes vital to managing risks and preventing shortages [31].
Reliable commodity price forecasts are essential for various stakeholders. Producers depend on price predictions to make informed planting and harvesting decisions. Importers and exporters utilize forecasts to negotiate trade agreements and manage price risks effectively. Traders and investors in financial markets rely heavily on accurate price forecasts for speculative investments and portfolio management, viewing commodities as valuable assets that can diversify risk and potentially offer stability. Governments and policymakers require precise forecasts for macroeconomic planning, managing essential goods’ supply-demand dynamics, making informed decisions about agricultural support programs, and developing sound economic policies. This broad reliance highlights the systemic importance of commodity price fluctuations; inaccuracies in forecasting can have widespread economic consequences, impacting everything from consumer purchasing power to national economic performance and international trade dynamics.
Despite its significance, commodity price forecasting remains notably challenging. Commodity markets exhibit inherent complexity and volatility, characterized by nonlinear relationships and nonstationary trends [42]. Prices respond dynamically to a broad range of interrelated factors including fundamental supply-demand variables (production levels, crop yields, and inventory), macroeconomic indicators (inflation rates, industrial output, interest rates), financial market influences (currency exchange rates, investor behavior, volatility measures), geopolitical incidents, trade disruptions, and environmental variables like extreme weather events and climate change, especially affecting agricultural commodities. Furthermore, these influencing factors constantly evolve. Increased financial market involvement in commodities, frequent global crises (economic downturns, pandemics), erratic climate patterns, and shifting geopolitical conditions all add further complexity and unpredictability. The growing intricacy of these factors intensifies volatility, rendering forecasting increasingly challenging and demanding adaptive models capable of dynamically integrating new information and capturing evolving relationships. Consequently, traditional forecasting methods often fall short, necessitating exploration into more advanced modeling approaches.
@@ 2.2 Traditional Econometric Approaches: Foundations and Limitations
The quantitative forecasting of commodity prices has historically been dominated by traditional econometric and time series models. These methods provided the foundational tools for analyzing price dynamics and generating predictions based on historical data. Prominent among these are linear regression models, Autoregressive Integrated Moving Average (ARIMA) models, Vector Autoregression (VAR) models, Vector Error Correction Models (VECM) for cointegrated series, and Random Walk models. ARIMA models, in particular, gained widespread popularity over several decades, valued for their well-defined statistical properties and the systematic identification procedure proposed by Box and Jenkins [13]. Studies have demonstrated ARIMA’s utility in specific contexts, such as forecasting lead and zinc prices or certain agricultural commodity prices under specific conditions [68].
Recognizing the pervasive issue of time-varying volatility (i.e., periods of high volatility followed by periods of relative calm, known as volatility clustering) in financial and commodity markets, Generalized Autoregressive Conditional Heteroskedasticity (GARCH) models were developed [12, 29]. GARCH and its numerous extensions specifically aim to model and forecast the conditional variance of price returns. This focus on volatility acknowledges that risk assessment and management are critical aspects of commodity market analysis.
Despite their established role, traditional econometric models face significant limitations when applied to the complexities of modern commodity markets, primarily their reliance on restrictive assumptions like linearity and stationarity [90]. Commodity price series frequently exhibit strong nonlinearity and structural breaks [76], violating these core assumptions. While some studies report adequate performance for traditional models in specific scenarios [10], a growing body of evidence highlights their inadequacy for many commodity prices. The recognized shortcomings of these models served as a catalyst for exploring more flexible modeling paradigms [77, 56].
@@ 2.3 Rise of Machine Learning in Commodity Forecasting
In response to the limitations of traditional models, the field has seen a significant shift towards machine learning (ML) and artificial intelligence (AI) techniques [19, 85]. ML models can model complex, nonlinear relationships directly from data without restrictive assumptions, often leading to improved forecasting accuracy compared to traditional benchmarks [69]. ML also facilitates analysis of diverse datasets, including unstructured data via NLP [99].
This reflects a potential paradigm shift from theory-driven econometrics to data-driven ML [16]. However, ML models, especially deep learning, often require large datasets [37], which can be a constraint in some commodity markets.
@@ 2.4 Key Machine Learning Techniques Applied
A spectrum of ML algorithms has been applied, as summarized in Table 1.
@@
Support Vector Machine (SVM) / Support Vector Regression (SVR) Finds optimal hyperplane to separate/regress data, using kernel tricks for non-linearity. Handles non-linearity effectively, robust baseline performance, good generalization. Sensitive to parameter/kernel choice, can be computationally intensive for large datasets.
Tree-based Ensembles (e.g., Random Forest (RF), Gradient Boosting) Combines multiple decision trees to improve prediction accuracy and robustness. Handles non-linearity, captures interactions, provides feature importance, robust to outliers. Can be prone to overfitting if not properly tuned, less interpretable than single trees.
@@
Long Short-Term Memory (LSTM) Type of RNN with gating mechanisms to control information flow and capture long-range dependencies. Excels at modeling sequential data and long-term dependencies. Computationally expensive, requires significant data, complex architecture.
Gated Recurrent Unit (GRU) Simplified version of LSTM with fewer gates. Captures long-term dependencies like LSTM, potentially more computationally efficient. Performance relative to LSTM is context-dependent.
Transformer Uses self-attention mechanisms to weigh input sequence importance, parallelizable. Excellent at capturing long-range dependencies. Computationally very expensive, requires large datasets.
@@
Support Vector Regression (SVR) [27] and tree-based ensembles like Random Forests (RF) [15] and XGBoost [20] are foundational. SVR handles nonlinearities well [93, 50, 74, 73]. Tree-based methods are robust and can model interactions [92].
Deep learning, starting with ANNs/MLPs [84, 102], evolved to Recurrent Neural Networks (RNNs) [28] for sequential data. LSTMs [46] specifically address RNNs’ vanishing gradient problem [8], making them effective for long-term dependencies in commodity forecasting [30, 103, 88]. Architectural variants like bidirectional LSTMs [86] and CNN-LSTM hybrids [53, 65] exist. GRUs [21] offer a simpler alternative, with performance varying by context [100, 22]. CNNs [58, 7] also find use, often in hybrids. The choice often involves trade-offs between complexity, data needs, and interpretability [44].
Transformer models [95], with their self-attention mechanism, represent a newer frontier, excelling at long-range dependencies [104, 62, 61, 64]. They are often combined with decomposition techniques in hybrid models [97]. The development of very large models (LLMs) also influences this space [17, 98].
Hybrid models combining different techniques are common [101, 52]. A prevalent strategy is ”decomposition and ensemble” [57], using methods like Wavelet Transforms [24], VMD [26], EMD [47], SSA [36], or STL [23] to simplify the series before forecasting each component [6, 40, 43, 60]. Explainable AI (XAI) is also being integrated [75].
Performance evaluation uses metrics like RMSE, MAE, and directional accuracy [48], sometimes with statistical tests like Diebold-Mariano [25]. ML/DL models often outperform traditional ones, but rankings vary by context [91, 49].
@@ 3.1 Problem Formulation
We consider a set of $d$ market features (e.g., macroeconomic, supply/demand, technical indicators) at time $t$, denoted by $\mathbf{x}_{t}\in\mathbb{R}^{d}$. The model input consists of a look-back window of $W$ historical observations:
$X=X_{t-W+1:t}=(\mathbf{x}_{t-W+1},\ldots,\mathbf{x}_{t})\in\mathbb{R}^{W\times d}$
Our objective is to predict the commodity futures price $p_{t}$ at $N$ distinct future horizons, where $\tau_{1}<\tau_{2}<\ldots<\tau_{N}$. The corresponding target vector is:
$Y=Y_{t}=(p_{t+\tau_{1}},p_{t+\tau_{2}},\ldots,p_{t+\tau_{N}})^{\top}\in\mathbb{R}^{N}$
The RSAE model aims to learn a mapping from $X$ to $Y$ via an intermediate sparse latent variable $\mathbf{z}\in\mathbb{R}^{m}$, where $m$ is the dimensionality of the latent space and $m\ll W\times d$. This latent variable $\mathbf{z}$ is intended to capture the essential underlying market state relevant for forecasting future price evolution.
@@ 3.2 Model Architecture
The RSAE architecture comprises three distinct neural network modules interacting within an energy-based learning paradigm. (A conceptual diagram of the architecture is presented in Figure 1.)
Figure 1: Conceptual architecture of the Regularized Sparse Autoencoder (RSAE) framework, illustrating the interaction between the History Predictor (Pred), Decoder (Dec), and Encoder (Enc) modules, along with the inference process for the sparse latent variable $\mathbf{z}$.
@@
    History Predictor (Pred): This module, parameterized by $\theta_{\text{Pred}}$, processes the historical input window $X$ to generate a condensed summary or context vector $\bm{h}\in\mathbb{R}^{h_{\text{dim}}}$. Given the sequential nature of $X$, this module typically utilizes recurrent layers (e.g., LSTMs [46], GRUs [21]) or temporal convolutional networks (1D CNNs) [7].
    $\bm{h}=\text{Pred}(X;\theta_{\text{Pred}})$
@@
    Decoder (Dec): This module, parameterized by $\theta_{\text{Dec}}$, takes both the history context $\bm{h}$ and a specific latent variable $\bm{z}\in\mathbb{R}^{m}$ as input to predict the multi-horizon future price vector $\tilde{Y}\in\mathbb{R}^{N}$.
    $\tilde{Y}=\text{Dec}(\bm{z},\bm{h};\theta_{\text{Dec}})$
    The latent variable $\bm{z}$ acts as a representation of the underlying factors modulating the future trajectory predicted from the historical context $\bm{h}$.
@@
    Encoder (Enc): This module, parameterized by $\theta_{\text{Enc}}$, performs amortized inference [33, 55]. Its role is to learn to directly predict the optimal latent variable $\bm{z}^{*}$ (derived via the inference process detailed below) given the actual future outcome $Y$ and the history context $\bm{h}$.
    $\hat{\bm{z}}=\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})$
    where $\hat{\bm{z}}$ is the Encoder’s estimate of $\bm{z}^{*}$.
@@ 3.3 Latent Variable Framework: Energy Function and Inference
Central to the RSAE is the determination of the latent variable $\bm{z}$. We define an energy function $E(Y,\bm{h},\bm{z})$ that quantifies the compatibility between a latent code $\bm{z}$, the historical context $\bm{h}$, and the future outcome $Y$. Low energy corresponds to a sparse latent code $\bm{z}$ that enables the Decoder to accurately predict $Y$ from $\bm{h}$ [59].
Energy Function
$E(Y,\bm{h},\bm{z})=C(Y,\text{Dec}(\bm{z},\bm{h};\theta_{\text{Dec}}))+\lambda R(\bm{z})$
where:
  • $C(Y,\tilde{Y})$ is the prediction cost, measuring the error between the true future $Y$ and the prediction $\tilde{Y}=\text{Dec}(\bm{z},\bm{h})$. We typically use Mean Squared Error (MSE) [11]:
    $C(Y,\tilde{Y})=\frac{1}{N}\left\|Y-\tilde{Y}\right\|_{2}^{2}$
  • $R(\bm{z})$ is the L1-norm sparsity regularizer [94]:
    $R(\bm{z})=\|\bm{z}\|_{1}=\sum_{k=1}^{m}|z_{k}|$
  • $\lambda>0$ is the hyperparameter controlling the strength of the sparsity penalty.
Optimal Latent Variable ($\bm{z}^{*}$)
For a given input pair $(X,Y)$ yielding $(\bm{h},Y)$, the optimal latent variable $\bm{z}^{*}$ is the value of $\bm{z}$ that minimizes the energy function:
$\bm{z}^{*}=\arg\min_{\bm{z}\in\mathbb{R}^{m}}E(Y,\bm{h},\bm{z})$
This minimization problem is typically non-convex if the decoder is non-linear.
Inference via Optimization
Finding $\bm{z}^{*}$ involves solving this optimization problem. We employ an iterative approach using gradient descent (or a subgradient method due to the L1 term) for a fixed number of steps $K$. This constitutes the explicit inference stage within the training loop:

### v2

@@ Abstract
Commodity futures price volatility creates significant economic challenges, necessitating accurate multi-horizon forecasting. Predicting these prices is complicated by diverse interacting factors (macroeconomic, supply/demand, geopolitical). Current models often lack transparency, limiting strategic use. This paper presents a Regularized Sparse Autoencoder (RSAE), a deep learning framework for simultaneous multi-horizon commodity futures prediction and discovery of interpretable latent market drivers. The RSAE forecasts prices at multiple horizons (e.g., 1-day, 1-week, 1-month) using multivariate time series. A key L1 regularization ($\|\mathbf{z}\|_{1}$) on its latent vector $\mathbf{z}$ enforces sparsity, promoting parsimonious explanations of market dynamics through learned factors representing underlying drivers (e.g., demand shifts, supply shocks). Drawing from energy-based models and sparse coding, the RSAE optimizes predictive accuracy while learning sparse representations. Evaluated on historical Copper and Crude Oil futures data with numerous indicators, our findings suggest the RSAE offers competitive multi-horizon forecasting accuracy and data-driven insights into price dynamics via its interpretable latent space, a notable advantage over traditional black-box approaches.
@@ 2.1 Introduction to Commodity Price Forecasting: Significance and Challenges
Accurate forecasting of commodity prices is crucial, given their extensive impact on economies and daily life [1]. Variations in prices of essential goods affect economic activities and individual well-being [36]. Macroeconomic stability of commodity-dependent countries is closely linked to price fluctuations [19], and predicting staple commodity prices is vital for food security [32]. Reliable forecasts are essential for producers, traders, investors, and policymakers.
Commodity price forecasting is notably challenging due to inherent market complexity, volatility, nonlinear relationships, and nonstationary trends [43]. Prices respond to numerous interrelated factors including supply/demand fundamentals, macroeconomic indicators, financial market influences, geopolitical events, and environmental variables. These factors evolve, with increased financialization, global crises, and climate change adding unpredictability, often rendering traditional methods inadequate.
@@ 2.2 Traditional Econometric Approaches: Foundations and Limitations
Quantitative commodity forecasting historically relied on models like ARIMA, VAR, and VECM. Box-Jenkins ARIMA models [14] gained popularity for their statistical properties [70]. GARCH models [13, 30] were developed to address time-varying volatility. However, these models struggle with the nonlinearities and structural breaks common in commodity prices [93, 80], often underperforming [11]. Their limitations spurred exploration of more flexible paradigms [81, 58].
@@ 2.3 Rise of Machine Learning in Commodity Forecasting
ML and AI techniques have become prominent alternatives [20, 88]. ML models inherently handle complex, nonlinear relationships without restrictive assumptions, often improving accuracy over traditional methods [71]. They can also integrate diverse data, including unstructured text via NLP [102]. This data-driven approach [17] contrasts with theory-driven econometrics but typically requires large datasets, a potential constraint in some markets [38].
@@ 2.4 Key Machine Learning Techniques Applied
A spectrum of ML algorithms has been applied (Table 1).
@@
Support Vector Machine (SVM) / Support Vector Regression (SVR) Finds optimal hyperplane to separate/regress data, using kernel tricks for non-linearity. Handles non-linearity effectively, robust baseline performance, good generalization. Sensitive to parameter/kernel choice, computationally intensive for large datasets.
Tree-based Ensembles (e.g., Random Forest (RF), Gradient Boosting) Combines multiple decision trees to improve prediction accuracy and robustness. Handles non-linearity, captures interactions, provides feature importance, robust to outliers. Prone to overfitting if not properly tuned, less interpretable than single trees.
@@
Long Short-Term Memory (LSTM) Type of RNN with specialized gating mechanisms (input, forget, output gates) to control information flow and capture long-range dependencies in sequential data. Excels at modeling sequential data and long-term dependencies, mitigating vanishing gradient problem of simple RNNs. Computationally expensive, requires significant data, complex architecture demanding careful tuning.
Gated Recurrent Unit (GRU) Simplified LSTM version with fewer gates (reset and update gates), also for long-range dependencies. Captures long-term dependencies like LSTM, potentially more computationally efficient due to fewer parameters. Performance relative to LSTM is context-dependent, may not always match LSTM’s capabilities on complex tasks.
Transformer Uses self-attention mechanisms to weigh input sequence importance globally, allowing parallel processing of sequence elements. Excellent at capturing very long-range dependencies, often outperforming RNNs/LSTMs on such tasks; parallelizable. Computationally very expensive (especially for long sequences), requires large datasets for effective training; newer in time series.
@@
Foundational models include SVR [28], noted for handling non-linearities [96, 52, 77, 76], and tree-based ensembles like RF [16] and XGBoost [21], valued for robustness and interaction modeling [95].
Deep learning (DL) advanced from ANNs/MLPs [87, 105] to RNNs [29]. LSTMs [48] and GRUs [22] became key for sequential data due to their ability to learn long-term dependencies by addressing the vanishing gradient issue [8]. They are widely used in commodity forecasting [31, 106, 91, 103, 23], with variants like bidirectional LSTMs [89] and CNN-LSTM hybrids [55, 67, 60, 7]. The choice involves balancing complexity, data needs, and interpretability [46].
Transformer models [98], leveraging self-attention, represent a significant advancement for long-range dependency capture [64, 63, 66]. Recent architectures like PatchTST [74] and FEDformer [107] highlight ongoing progress. Transformers are often integrated into hybrid models, frequently with decomposition methods [100]. Large Language Models (LLMs) are also influencing financial forecasting research [18, 101].
Hybrid models are prevalent, often combining traditional methods with ML, or using a ”decomposition and ensemble” strategy [104, 54, 59]. Decomposition techniques (e.g., Wavelets [25], VMD [27], EMD [49], SSA [37], STL [24]) simplify series for component-wise forecasting [6, 41, 45, 62]. Integration of XAI is also an emerging trend [78].
Performance is typically assessed using RMSE, MAE, and directional accuracy [50]. Statistical tests like Diebold-Mariano [26] are sometimes used for comparing model accuracy. While ML/DL models often show superior performance, rankings are context-dependent [94, 51].
@@ 3.1 Problem Formulation
We consider a set of $d$ market features at time $t$, $\mathbf{x}_{t}\in\mathbb{R}^{d}$. The model input is a look-back window $X=X_{t-W+1:t}=(\mathbf{x}_{t-W+1},\ldots,\mathbf{x}_{t})\in\mathbb{R}^{W\times d}$. Our objective is to predict the futures price $p_{t}$ at $N$ horizons $\tau_{1}<\ldots<\tau_{N}$, with target $Y=Y_{t}=(p_{t+\tau_{1}},\ldots,p_{t+\tau_{N}})^{\top}\in\mathbb{R}^{N}$. The RSAE learns a mapping from $X$ to $Y$ via a sparse latent variable $\mathbf{z}\in\mathbb{R}^{m}$ ($m\ll W\times d$).
@@ 3.2 Model Architecture
The RSAE comprises three neural network modules (Figure 1).
Figure 1: Conceptual architecture of the Regularized Sparse Autoencoder (RSAE) framework.
@@
    History Predictor (Pred): $\bm{h}=\text{Pred}(X;\theta_{\text{Pred}})$, maps $X$ to a context vector $\bm{h}\in\mathbb{R}^{h_{\text{dim}}}$, using LSTMs [48], GRUs [22], or 1D CNNs [7].
@@
    Decoder (Dec): $\tilde{Y}=\text{Dec}(\bm{z},\bm{h};\theta_{\text{Dec}})$, predicts future prices $\tilde{Y}\in\mathbb{R}^{N}$ from $\bm{z}$ and $\bm{h}$.
@@
    Encoder (Enc) (for training): $\hat{\bm{z}}=\text{Enc}(Y,\bm{h};\theta_{\text{Enc}})$, performs amortized inference of $\bm{z}^{*}$ from actual future $Y$ and $\bm{h}$ [34, 57]. A separate encoder, $\text{Enc}_{\text{XtoZ}}(X;\theta_{\text{XtoZ}})$, is trained to predict $\bm{z}^{*}$ from $X$ alone for test-time forecasting. $\text{Enc}_{\text{XtoZ}}$ is trained with a mean-squared error objective to regress the K-step optimal z*; its parameters are not updated through the prediction loss.
@@ 3.3 Latent Variable Framework: Energy Function and Inference
We define an energy function $E(Y,\bm{h},\bm{z})=C(Y,\text{Dec}(\bm{z},\bm{h};\theta_{\text{Dec}}))+\lambda R(\bm{z})$ [61]. $C(Y,\tilde{Y})=\frac{1}{N}\|Y-\tilde{Y}\|_{2}^{2}$ is the MSE prediction cost [12]. $R(\bm{z})=\|\bm{z}\|_{1}$ is the L1-norm sparsity regularizer [97]. The optimal latent variable $\bm{z}^{*}=\arg\min_{\bm{z}}E(Y,\bm{h},\bm{z})$ is found via $K$-step (sub)gradient descent [15].
3.3.1 Selection and Convergence of Inference Steps (K)
The number of iterative steps, $K$, in the inference process for $\bm{z}^{*}$ is a hyperparameter balancing thoroughness of energy minimization against computational cost. To determine an appropriate value for $K$, preliminary experiments were conducted on the validation set. This involved tracking the average energy function value, $E(Y,\bm{h},\bm{z})$, over a range of potential iterations (e.g., $k=1$ to $k=15$) for multiple representative batches of validation samples.
The observed convergence behavior is depicted in Figure 2. The mean energy (averaged across validation batches) typically exhibited a sharp decrease during the initial iterations (approximately $k=1$ to $k=7$). Beyond this initial phase, the rate of energy reduction substantially diminished, with the curve beginning to plateau around $k=8$ to $k=10$. While formal convergence to a global minimum is not guaranteed for non-convex energy landscapes, the iterative process aims to find a suitable local minimum that provides a sparse and predictive latent code. Based on the empirical observation that iterations beyond $K=10$ provided only marginal decreases in the objective energy at an increased computational cost, $K=10$ was selected for all subsequent experiments in this study. A more extensive analysis of convergence properties under varying conditions remains an avenue for future work.
Figure 2: Mean objective energy function value $E(Y,\bm{h},\bm{z})$ versus the number of inference iterations ($k$), averaged over representative validation set batches. The plot demonstrates typical convergence behavior, with energy decreasing rapidly initially before plateauing. The selected $K=10$ (dashed red line) represents a point of diminishing returns for energy reduction.
3.4 Test-Time Inference
At test time, future $Y$ is unknown. We train an additional inference encoder, $\text{Enc}_{\text{XtoZ}}(X;\theta_{\text{XtoZ}})$, supervised by the $\bm{z}^{*}$ values obtained during the training phase (from the $K$-step optimization using known $Y_{\text{train}}$). Thus, for a new test input $X_{\text{test}}$, the latent code is obtained directly: $\bm{z}_{\text{test}}=\text{Enc}_{\text{XtoZ}}(X_{\text{test}};\theta_{\text{XtoZ}})$. This amortized inference from $X$ alone is used for generating forecasts.
3.5 Training Objective and Parameter Updates
The primary training objective minimizes the prediction loss using the $K$-step inferred $\bm{z}^{*}$: $L_{\text{Pred}}^{*}=C(Y,\text{Dec}(\bm{z}^{*},\bm{h};\theta_{\text{Dec}}))$. Parameters $\theta_{\text{Pred}}$ and $\theta_{\text{Dec}}$ are updated based on gradients from $L_{\text{Pred}}^{*}$. If the training-time Encoder $\text{Enc}(Y,\bm{h})$ is used, it is trained with an inference loss $L_{\text{Inf}}=\|\bm{z}^{*}-\text{Enc}(Y,\bm{h})\|_{2}^{2}$. The test-time encoder $\text{Enc}_{\text{XtoZ}}(X)$ is trained separately to predict $\bm{z}^{*}$ from $X$.
3.6 Overall Training Algorithm